# Understanding the approach to thermalization in a non-Abelian gauge theory

### v1

@@ 6 Outlook
In this letter, we provide a detailed understanding of the dynamical properties of soft (magnetic) modes of non-Abelian SU(2) gauge theory as a function of temperature as well as in a particular non-thermal attractor state. These magnetic modes interact non-perturbatively even at asymptotically high temperatures [40] and influence dynamical properties like the sphaleron transition rate [25, 41], and other transport properties of the non-Abelian plasma [42]. We show here that these modes play an important role in the process of thermalization in gauge theories. In a high temperature plasma, where a clean separation between the hard (ultraviolet momentum) modes and these soft modes is possible, resulting in an effective stochastic description of the latter [25], we demonstrate the chaotic behavior inherent in the system by extracting the spectra of positive Lyapunov exponents. This allows us to measure the Kolmogorov-Sinai entropy rate which determines how fast the entropy flows into the system eventually attaining a maximum. We have used this rate to measure a thermalization time $\sim 0.7$-$0.5$ fm/c to reach to a thermal state at $450$-$600$ MeV starting from a particular non-thermal state of soft gluons which shows a particular self-similar behavior [14]. The only assumption that goes into our calculation is the conservation of energy density of the soft modes during the entire evolution of the system. Interestingly our estimate of the thermalization time is consistent with the early switch-on time of hydrodynamics to describe flow observables in heavy-ion collisions [43] but is smaller than estimates ($\gtrsim 2.5$ fm/c) based on perturbative scattering process [44].
@@ 7 Acknowledgements
We are grateful to Sumilan Banerjee, Rajiv Gavai and Soeren Schlichting for discussions related to this work. The authors acknowledge the computing time provided by the High Performance Computing Center at the Institute of Mathematical Sciences.

### v2

@@ 6 Outlook
In this letter, we provide a detailed understanding of the dynamical properties of soft (magnetic) modes of non-Abelian SU(2) gauge theory as a function of temperature as well as in a particular non-thermal attractor state. These magnetic modes interact non-perturbatively even at asymptotically high temperatures [45] and influence dynamical properties like the sphaleron transition rate [28, 46], and other transport properties of the non-Abelian plasma [47]. We show here that these modes play an important role in the process of thermalization in gauge theories as well. In a high temperature SU(2) plasma, where a clean separation between the hard (ultraviolet momentum) modes and these soft modes is possible, resulting in an effective stochastic description of the latter [28], we demonstrate the chaotic behavior inherent in the system by extracting the spectra of positive Lyapunov exponents. This allows us to measure the Kolmogorov-Sinai entropy rate which determines how fast the entropy flows into the system eventually attaining a maximum. We have used this rate to measure a thermalization time $\sim 0.7$-$0.5$ fm/c to reach to a thermal state at $450$-$600$ MeV starting from a particular non-thermal state of soft gluons which shows a particular self-similar behavior [14]. The only assumption that goes into our calculation is the conservation of energy density of the soft modes during the entire evolution of the system. Interestingly our estimate of the thermalization time in this simple non-expanding system of overoccupied soft gluon modes is consistent with the early switch-on time for hydrodynamics necessary to describe flow observables in a heavy-ion collision event [48] which is described by a glasma initial condition in an expanding spacetime geometry. However our result for the thermalization time is smaller than typical estimates ($\gtrsim 2.5$ fm/c) based on perturbative scattering process [49].
@@ 7 Acknowledgements
We are grateful to Sumilan Banerjee, Rajiv Gavai, Sören Schlichting and Ravi Shanker for discussions related to this work. We also thank the anonymous Referee for several suggested improvements. The authors acknowledge the computing time provided by the High Performance Computing Center at the Institute of Mathematical Sciences.